# Remasking Discrete Diffusion Models with Inference-Time Scaling

### v1

@@ Appendix E Additional Experimental Results
In Table 3, we report QM9 dataset validation set perplexities for models trained with either the MDLM NELBO from (3) or the ReMDM NELBO from (11), with $\sigma_{t}=\min\{\eta_{cap},(1-\alpha_{s})/\alpha_{t}\}$, for some $\eta_{cap}\in(0,1]$. We follow the same model architecture and training settings defined in Schiff et al. (2024) for this dataset. Overall, we find that validation perplexities for this dataset are fairly consistent across models trained with either objective.
Table 3: Training with ReMDM and MDLM NELBO objectives leads to similar validation set perplexity values (for the QM9 dataset). In the top row, the model is trained using the continuous time formulation of the MDLM objective from Sahoo et al. (2024). In the bottom two rows, models are trained with discrete-time objectives. The first column, $\sigma_{t}=0$, corresponds to the MDLM objective from (3) and the columns with varying $\eta_{cap}$ correspond to training with the ReMDM objective from (11) with $\sigma_{t}=\min\{\eta_{cap},(1-\alpha_{s})/\alpha_{t}\}$. We find that results are comparable across training settings.
@@ E.2 OpenWebText
In Figure 6, we plot the remasking probability schedules of ReMDM and DFM correctors in the experiments reported in Table 1. Similarly to Figure 5, the DFM schedules show a spike at first and a long tail afterward. Here, the best performing $\beta_{t}=At^{0.25}(1-t)^{0.25}$ schedule reported in Gat et al. (2024) is used. We also adjust the width of the time steps accordingly, as reported in Gat et al. (2024).
E.2.2 Results of Models without Nucleus Sampling
In Table 4, we present the sample quality of different models without nucleus sampling. Consistent with the results in Zheng et al. (2024), when using double-precision floating-point numbers during sampling, discrete diffusion models produce poor quality sequences, as reflected by the relatively higher generative perplexity values compared to Table 1. These results underscore the importance of nucleus sampling for discrete diffusion models in text generation.
Table 4: Sample quality of absorbing state discrete diffusion models without nucleus sampling.
Method MAUVE ($\uparrow$) Gen PPL. ($\downarrow$) Entropy ($\uparrow$)
AR (T=1024) 0.407 35.0 5.58
T=1024 T=2048 T=4096 T=1024 T=2048 T=4096 T=1024 T=2048 T=4096
SEDD 0.008 0.008 0.009 104.7 103.2 102.5 5.62 5.61 5.61
MDLM 0.009 0.008 0.009 104.8 104.4 104.1 5.63 5.63 5.63
MDLM+FB 0.009 0.011 0.010 102.6 101.9 100.9 5.65 5.65 5.64
MDLM+DFM 0.008 0.010 0.010 125.9 119.1 115.7 5.70 5.69 5.69
ReMDM 0.007 0.007 0.007 173.0 256.2 370.6 5.76 5.81 5.74
T=128 T=256 T=512 T=128 T=256 T=512 T=128 T=256 T=512
SEDD 0.007 0.007 0.008 119.2 110.1 107.2 5.65 5.63 5.62
MDLM 0.007 0.008 0.008 121.4 111.8 107.2 5.67 5.65 5.64
MDLM+FB 0.007 0.009 0.010 112.8 106.4 104.4 5.67 5.66 5.65
MDLM+DFM 0.005 0.006 0.007 360.7 189.5 143.5 5.85 5.77 5.73
ReMDM 0.006 0.007 0.006 160.5 176.8 226.9 5.73 5.75 5.79
E.2.3 Comparing Different ReMDM Schedules
In Table 5, we report the OpenWebText unconditional generation results for various ReMDM schedules. For ReMDM-conf, we set $\sigma_{t}$ proportional to the upper bound $\sigma_{max}$. In the case of inference-time scaling ($T\geq 1024$), ReMDM-loop performs the best while in the case of faster sampling ($T<1024$), ReMDM-cap is found to perform better than others. In both scenarios, ReMDM-rescale and ReMDM-conf do not perform as well as the other two schedules.
Table 5: Comparing sample quality on OWT for various ReMDM schedules. indicates nucleus sampling. For each $T$, the best MAUVE score is bolded.
@@ E.2.4 Tuning $\eta_{cap}$ / $\eta_{rescale}$
In Figure 7, we plot MAUVE scores when using the ReMDM-cap schedule with different $\eta_{cap}$ values and varying number of decoding steps $T$. As shown in Figure 7(a), in the case of inference-time scaling ($T\geq 1024$), the MAUVE scores at $T=4096$ tend to decrease as $\eta_{cap}$ increases while the MAUVE scores at $T=1024$ tend to increase with larger $\eta_{cap}$. We attribute this to the perplexity-entropy trade-off. In particular, when $\eta_{cap}$ increases, the probability of remasking increases, improving perplexity, but also harming diversity. In the case of $T=4096$, the models can generate higher quality sentences and the reduction in diversity outweighs marginal improvements in generative perplexity. For $T=1024$, the samples are of relatively poorer quality and the improvement in perplexity plays a major role in driving up the MAUVE score. To achieve the best balance, we choose $\eta_{cap}=0.008$ for the setting of inference-time scaling.
For faster sampling ($T<1024$), in Figure 7(b), with the exception of $\eta=0.045$ we see a positive trend between increasing $\eta_{cap}$ and improving MAUVE scores. In Tables 1 and 5, we report results using $\eta=0.040$.
We also performed similar studies for schedules that combine ReMDM-cap with the ReMDM-loop strategy (Figure 8) and for ReMDM-rescale schedules (Figure 9). For both of these settings, we observe trends for increasing $\eta_{cap}$ / $\eta_{rescale}$ that are similar to those described above. In Table 5, we report the following choice of hyperparameter for each of these settings. For ReMDM-loop combined with ReMDM-cap, in the inference-time scaling experiments ($T\geq 1024$), we report results using $\eta_{cap}=0.020$. For the faster sampling experiments ($T<1024$), we report results for $\eta_{cap}=0.055$. For the ReMDM-rescale inference-time scaling experiments, we report results with $\eta_{rescale}=0.015$ and for the faster sampling experiments, we use $\eta_{rescale}=0.045$ in Table 1.
@@ E.3 ImageNet
In Table 6, we present the full results for varying the sampling temperature $\tau$ and the various ReMDM schedules and the their corresponding hyperparameters ($\eta_{cap}$ for ReMDM-cap. and $\eta_{rescale}$ for ReMDM-rescale). As noted in Section 5.1.2, while both MDLM and ReMDM benefit from using a softmax temperature of $\tau=0.8,$ the MaskGiT sampler produces the best results with no temperature scaling (i.e., $\tau=1$). Of note, we use FID to determine which setting constitutes the ‘best’ results for each sampler. With reduced temperature, the entropy of the softmax is reduced leading to less diverse samples. This is reflected in the trade-off of the FID vs. IS metrics across sampling temperatures, with IS penalizing a lack of diversity less than FID (Heusel et al., 2017).
@@ E.4 QM9
In Figures 11-26, we present the results from the extensive hyperparameter search we conducted across both properties, ring count (Figures 11-18) and QED (Figures 19-26), and guidance mechanisms D-CFG (Figures 11-14, 19-22) and D-CBG (Figures 15-18, 23-26).
@@ Appendix G Assets
ImageNet (Deng et al., 2009) ImageNet License
OpenWebText (Gokaslan & Cohen, 2019) Creative Commons CC0 license (“no rights reserved”)
QM9 (Ruddigkeit et al., 2012; Ramakrishnan et al., 2014) N/A
@@
Guided Diffusion (Dhariwal & Nichol, 2021) MIT
HuggingFace (Wolf et al., 2019) Apache 2.0
Hydra (Yadan, 2019) MIT
Jax (Bradbury et al., 2018) Apache 2.0
NumPy (Harris et al., 2020) NumPy license
MaskGiT (Chang et al., 2022) Apache 2.0
Matplotlib (Hunter, 2007) Matplotib license
@@
Pandas (pandas development team, 2020) BSD 3-Clause “New” or “Revised”
PyTorch (Paszke et al., 2019) BSD-3 Clause
PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019) Apache 2.0
RDKit (Landrum et al., 2013) BSD 3-Clause “New” or “Revised”
Seaborn (Waskom, 2021) BSD 3-Clause “New” or “Revised”
@@
—endoftext— might be seen as a tough choice, it’s smart to do so. According to past reports, the SIM cards are only made available in the powerful manufacturer’s range of carrier variants. And though the ZTE does its own case with these phones, this variant will not be disclosed. We also know the color of the phone and what it also will cost remains to be seen. With the Carrington showdown, the new edition of the exclusive TVD features more pictures with Samsung. —endoftext— Goodes The Derby winner was named first-team of the Match on the BBC evening. He feels his position as manager at Southall last season was a great honour and understands the surarious delight in earning that he can’t complain about delivering an endless ovation. The Newcastle coach knows how describes how people like their approach to games. As his biggest fans, he has criticised Paul Cruy and traded barbs with Chris Froome. And sad to admit, he admits he did not get everything right. He just betrayed them in his intention and left an exasperation in others. I’ve finished telling players that we’d get that, and they continued that attitude to how we didn’t think we deserved the job ©Newcastle United Rovers manager Tony Wheeler All this sound of being needed for interviews in various places. “I’m quite pleased by the passion the fans have shown for me but believe me, I’m not the fan - I find the person out. It would be an extremely nice position to be in. Illuminate. We’d like to think, not to get a player based on Newcastle’s team when he does take charge. “My head coach says, ’Listen, it’s just too much to cry,’ Tony Wheeler told the BBC on Tuesday about the injury he played in the Northern League final at Paroles. “If the player is a part in the starting team, then I’m confident he would be put in there, but I also’ve finished telling players we thought we’d get the job. We had good options to go in the squad. “We had, obviously, the people that were always willing for us to get the best player we could because as coaches we felt we had the best players, and we felt that no one had a choice. “And the only thing [I] was just going on to say was that we were aware beforehand that it was an unusual situation for us to hook him up with the potential of [Meagan] Forster. “We like his game and he likes ours, and if he’s a good striker, he can go out there and play both for the team and for ourselves. “Stephen [Biggish] Jackson, who’s worked with here, obviously respects me with great respect and likes his opinions, but I think [Tony Wheeler] can say that it’s satisfactory. “We’ve got lots of many, and them combine them with Van Ferdinand. All three people we’ve been approached to, they need to be taken. “Could something be done to them? Not be done.” United are certainly prepared to lose their star man for Tuesday’s Rovers clash, with Borby to take charge of their trip to Dunferm on Saturday. Although they have won five of their six games, they sit sixth in the Northern League, which includes the Parkhead and the two assistant charges but on Tuesday their only goal of the season is worthy of their reward. United have more goals than Palace, scoring both goals this season despite the best form of Johnson as new coach Winger Darius Boyd joined Derby after replacing Marlon Wickham in June and recently, he stayed for a year on loan until he joined Newcastle in defiance of Tony Wheeler. Tom Williams, a future United captain, defender and striker, told BBC One Sports: “Being in Rovers is great for them. “They develop their chemistry very well in the dugout and they try to stay in good relationships together. “They’re unabashed about that you can play with both of them now, so they’ve stayed loyal.” Brendan Rodgers, the current Newcastle United manager after Wheeler took over from Mick Jagger, and succeeded David Moyes, as the most paid Premier League manager in 1997 while the Scot was the highest paid Everton manager. Leicester manager Leicester Leicester is yet to make terms with the joys of the job and has become so indebted he offered to forgive Newcastle United’s owners for the cost of lost profits after buying a stake in the club earlier this year. —endoftext— Sometimes, I enjoy having colleagues “embarrassed” to compete in sports, especially it comes to being a one-richen employee’s career. I can’t own the jobs of the colleagues that pressure us from Chinese and Australian competitive athletes and even to compete in the United States. They seem arrogant —endoftext—
@@
Figure 27: Sample from MDLM and $T=4096$ diffusion steps. —endoftext— token is bolded to emphasize the boundary between paragraphs.
@@
—endoftext— a component to renovating the park, but no one wants to involve others in it before facing a backlash. “We are going to stand up and try to make a real statement on this,” Thomas told reporters at a rally Tuesday night. “I would much rather have that there be a consensus not just to make a decision, but to have a very real discussion, and to begin to build a movement.” The idea behind the proposal is a call to the public, council and the youth advisory board, which was approved by residents on the first day of the meeting. Thomas said the message will be important to Hamilton youth. “It’s going to be very much about a number of people,” he said. “There will be arguments on both sides. It is not going to be all about the type of person that you are or the committee member that you are with a particular number of people.” Thomas said that the plan is really about an initiative to help the community do both of those things. “This isn’t going to be a positive thing, as it should be, but it’s going to put people in position to make sure that our focus is not just about how to deal with poverty …and with homelessness. Our focus is on keeping the people in our community affordable and to create opportunities for their families. “We have —believe it or not —we have increased community involvement in our city,” Thomas said. “I think it’s obvious from all of the issues now that we’ll consider resource-generating development and how to devribute those resources.” The renovated park is one of the 14 parks identified in the city’s 2008 map of the park’s 2-acre site, which was opened up for a community arts festival. City staff started coming up with the project in late December, but the public has yet to come to terms. For their part, residents sign the petition only to receive two differing opinions. Although there are concerns that it would be an uphill battle that requires council approval, officials said Tuesday that the process could eventually take shape, with councillors likely to support it. Hamilton city hall spokesman John Little told reporters that the city would have to address a petition from the public if there is one on site. “We are not able to change the outcome because we are discussing it internally, so what we have has not made an official decision,” Little said. Thomas said staff wants to hear how the public feels about the project, and it must be incorporated in such a prompt fashion that the city is looking at more permanent changes. “We implemented a 30-year plan, about 10 years ago, and we will continue it this year,” Thomas said. “We want to once again really emphasize community involvement. I mean, as we say we take this kind of step, we don’t want to make separate efforts to change different things,” he said. Any plans would be discussed with community boards, Conley said. Hamilton city councillor Steve Cowen said if residents sign the petition, “actions have to be done, knowing when these changes are needed and that we will have the right options to move forward.” Cowen said two renovations along with suggestions for revenue set by the Hulbert Company, which operates the Gerrard Elementary school theatre, will be included on city documents and website. They would set aside \$2 million more than the result of Washtenaw Park Plaza renovation plans, sparked by the QEDA expansion that renovated the former Washtenaw Historic Restaurant and Bakery on Gerrard. The company has also purchased Gerrard and turned it into a New American Food, Inc. restaurant, and brought the business before the city council in late 2015. Hamilton city council has been having this conversation for more than a year, but it usually happens only now. “It’s really not, ’We need a meeting soon,’ why we need to do this,” Conley said. “It’s just a forum where people can voice their views and inform council, and within the next year, hopefully council will do the right thing to do.” With John Elliott: 905-526-2431 mmelliott@thespec.com —endoftext— Written by Steven B. Ollander In the last five years, we have seen huge financial crises around the world, and not just economic ones. We’ve also seen rapid changes in economic systems, in central and peripheral countries. All these systems are in the hands of banks, but there are other forces that have also been used to dump more money into other countries —endoftext—

### v2

@@ Appendix E Additional Experimental Results
In Table 4, we report QM9 dataset validation set perplexities for models trained with either the MDLM NELBO from (3) or the ReMDM NELBO from (3.2), with $\sigma_{t}=\min\{\eta_{cap},(1-\alpha_{s})/\alpha_{t}\}$, for some $\eta_{cap}\in(0,1]$. We follow the same model architecture and training settings defined in Schiff et al. [43] for this dataset. Overall, we find that validation perplexities for this dataset are fairly consistent across models trained with either objective.
Table 4: Training with ReMDM and MDLM NELBO objectives leads to similar validation set perplexity values (for the QM9 dataset). In the top row, the model is trained using the continuous time formulation of the MDLM objective from Sahoo et al. [38]. In the bottom two rows, models are trained with discrete-time objectives. The first column, $\sigma_{t}=0$, corresponds to the MDLM objective from (3) and the columns with varying $\eta_{cap}$ correspond to training with the ReMDM objective from (3.2) with $\sigma_{t}=\min\{\eta_{cap},(1-\alpha_{s})/\alpha_{t}\}$. We find that results are comparable across training settings.
@@ E.2 OpenWebText
In Figure 5, we plot the remasking probability schedules of ReMDM and DFM correctors in the experiments reported in Table 1. The DFM schedules show a spike at first and a long tail afterward. Here, the best performing $\beta_{t}=At^{0.25}(1-t)^{0.25}$ schedule reported in Gat et al. [13] is used. We also adjust the width of the time steps accordingly, as reported in Gat et al. [13].
E.2.2 Hacking Generative Perplexity
In Table 5, we present the MAUVE, Gen PPL., and entropy values of ReMDM-cap on OWT with different ReMDM hyperparameters $\eta$. As $\eta$ increases, Gen PPL. decreases drastically at the sacrifice of entropy/diversity. In other words, Gen PPL. is hacked by tuning hyperparameters. In contrast, MAUVE cannot be hacked in this way since it measures the balance between generative perplexity and diversity by definition.
Table 5: As $\eta$ increases, we are able to get arbitrarily good Gen PPL. at the sacrifice of entropy/diversity. In contrast, MAUVE cannot be hacked in this way. ReMDM-cap on OWT with $T=1024$ and $L=1024$ is used and for each experiment, we generate 5000 samples. For each metric, the best value is bolded.
Metric MAUVE ($\uparrow$) Gen PPL. ($\downarrow$) Entropy ($\uparrow$)
$\eta_{cap}=0.008$ 0.36 27.7 5.32
$\eta_{cap}=0.05$ 0.25 12.9 4.93
$\eta_{cap}=0.2$ 0.02 3.1 2.43
E.2.3 Comparing Different ReMDM Strtegies
In Table 6, we report the OpenWebText unconditional generation results for various ReMDM schedules. In the case of inference-time scaling ($T\geq L$), ReMDM-loop performs the best while in the case of faster sampling ($T<L$), ReMDM-cap is found to perform better than others. In both scenarios, ReMDM-rescale does not perform as well as the other two strategies.
Table 6: Comparing sample quality on OWT for various ReMDM strategies. indicates nucleus sampling. For each $T$, the best MAUVE score is bolded.
@@ E.2.4 Tuning $\eta_{cap}$ / $\eta_{rescale}$
In Figure 6, we plot MAUVE scores when using the ReMDM-cap schedule with different $\eta_{cap}$ values and varying number of decoding steps $T$. As shown in Figure 6(a), in the case of inference-time scaling ($T\geq L$), the MAUVE scores at $T=4096$ tend to decrease as $\eta_{cap}$ increases while the MAUVE scores at $T=1024$ tend to increase with larger $\eta_{cap}$. We attribute this to the perplexity-entropy trade-off. In particular, when $\eta_{cap}$ increases, the probability of remasking increases, improving perplexity, but also harming diversity. In the case of $T=4096$, the models can generate higher quality sentences and the reduction in diversity outweighs marginal improvements in generative perplexity. For $T=1024$, the samples are of relatively poorer quality and the improvement in perplexity plays a major role in driving up the MAUVE score. To achieve the best balance, we choose $\eta_{cap}=0.008$ for the setting of inference-time scaling.
For faster sampling ($T<L$), in Figure 6(b), with the exception of $\eta=0.045$ we see a positive trend between increasing $\eta_{cap}$ and improving MAUVE scores. In Tables 1 and 6, we report results using $\eta=0.040$.
We also performed similar studies for schedules that combine ReMDM-cap with the ReMDM-loop strategy (Figure 7) and for ReMDM-rescale schedules (Figure 8). For both of these settings, we observe trends for increasing $\eta_{cap}$ / $\eta_{rescale}$ that are similar to those described above. In Table 6, we report the following choice of hyperparameter for each of these settings. For ReMDM-loop combined with ReMDM-cap, in the inference-time scaling experiments ($T\geq L$), we report results using $\eta_{cap}=0.020$. For the faster sampling experiments ($T<L$), we report results for $\eta_{cap}=0.055$. For the ReMDM-rescale inference-time scaling experiments, we report results with $\eta_{rescale}=0.015$ and for the faster sampling experiments, we use $\eta_{rescale}=0.045$ in Table 1.
@@ E.3 ImageNet
In Table 7, we present the full results for varying the sampling temperature $\tau$ and the various ReMDM schedules and the their corresponding hyperparameters ($\eta_{cap}$ for ReMDM-cap. and $\eta_{rescale}$ for ReMDM-rescale). As noted in Section 5.1.2, while both MDLM and ReMDM benefit from using a softmax temperature of $\tau=0.8,$ the MaskGiT sampler produces the best results with no temperature scaling (i.e., $\tau=1$). Of note, we use FID to determine which setting constitutes the ‘best’ results for each sampler. With reduced temperature, the entropy of the softmax is reduced leading to less diverse samples. This is reflected in the trade-off of the FID vs. IS metrics across sampling temperatures, with IS penalizing a lack of diversity less than FID [16].
@@ E.4 QM9
In Figures 10-25, we present the results from the extensive hyperparameter search we conducted across both properties, ring count (Figures 10-17) and QED (Figures 18-25), and guidance mechanisms D-CFG (Figures 10-13, 18-21) and D-CBG (Figures 14-17, 22-25).
@@ E.4.4 Tuning $t_{switch}$ / Loop Parameters
E.4.5 Comparing to FB and DFM Correctors
In Table 8, we compare ReMDM to the other predictor-corrector samplers, namely FB [5] and DFM [13] in the context of maximizing ring count using conditional generation (i.e., $\gamma=1$ for D-CFG). ReMDM better trades-off novel sample generation and maximizing the property of interest.
Table 8: ReMDM better trades-off sample novelty and ring count maximization compared to other predictor-corrector methods.
Sampler Num Novel Novel Ring Count Mean
$T=32$
FB 268 4.35
DFM 341 4.57
ReMDM 273 4.64
$T=64$
FB 304 4.42
DFM 311 4.57
ReMDM 318 4.61
$T=128$
FB 295 4.53
DFM 236 4.52
ReMDM 348 4.53
@@ Appendix F Generated Samples
F.1 Generative Perplexity Hacking Example
We visualize a generative perplexity hacking example in Figure 26. In this example, ReMDM-cap on OWT with $\eta_{cap}=0.2$ and $T=1024$ generates a sequence filled with repetitive tokens without real semantics, but the generative perplexity is as low as 1.5.
F.2 OpenWebText
@@
F.3 LLaDA
We visualize examples of LLaDA answering Countdown in Figure 30, and LLaDA answering TruthfulQA questions in Figure 31. In both cases, we use few-shot evaluation. Particularly, we use 4-shot for Countdown and 6-shot for TruthfulQA. The few shot examples presented in the prompt are the same for all questions in one task.
@@ Appendix G Assets
ImageNet [8] ImageNet License
OpenWebText [14] Creative Commons CC0 license (“no rights reserved”)
QM9 [37, 36] N/A
TruthfulQA
[24] Apache-2.0
@@
Guided Diffusion [9] MIT
HuggingFace [56] Apache 2.0
Hydra [57] MIT
Jax [4] Apache 2.0
NumPy [15] NumPy license
MaskGiT [7] Apache 2.0
Matplotlib [20] Matplotib license
@@
Pandas [30] BSD 3-Clause “New" or “Revised"
PyTorch [31] BSD-3 Clause
PyTorch Lightning [11] Apache 2.0
RDKit [21] BSD 3-Clause “New" or “Revised"
Seaborn [54] BSD 3-Clause “New" or “Revised"
MAUVE [33, 34] GPLv3
Language Model Evaluation Harness [12] MIT
<|endoftext|>, my love, my love, my love, my love, my love, my love, oh, oh, oh, oh, oh, oh, oh, oh, my love, my love, my love, my love, oh, oh, oh, oh, oh, oh, oh, oh, oh, oh, oh, oh, oh, oh, oh, oh, my love, my love, my love, my love, my love, my love, my love, hate, hate, hate, hate, hate, hate, hate, hate, hate, hate, hate, hate, hate, hate, hate, hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate HHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHH HHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHH HHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHH hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama Mama hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate hate<|endoftext|>
Figure 26: Example of generative perplexity hacking. The generated sequence by ReMDM-cap on OWT with $\eta_{cap}=0.2$ and $T=1024$ features repetitive words without real semantics. Gen PPL. is 1.5 and entropy is 1.04. |endoftext| token is bolded to emphasize the boundary between paragraphs.
@@
|endoftext| might be seen as a tough choice, it’s smart to do so. According to past reports, the SIM cards are only made available in the powerful manufacturer’s range of carrier variants. And though the ZTE does its own case with these phones, this variant will not be disclosed. We also know the color of the phone and what it also will cost remains to be seen. With the Carrington showdown, the new edition of the exclusive TVD features more pictures with Samsung. |endoftext| Goodes The Derby winner was named first-team of the Match on the BBC evening. He feels his position as manager at Southall last season was a great honour and understands the surarious delight in earning that he can’t complain about delivering an endless ovation. The Newcastle coach knows how describes how people like their approach to games. As his biggest fans, he has criticised Paul Cruy and traded barbs with Chris Froome. And sad to admit, he admits he did not get everything right. He just betrayed them in his intention and left an exasperation in others. I’ve finished telling players that we’d get that, and they continued that attitude to how we didn’t think we deserved the job ©Newcastle United Rovers manager Tony Wheeler All this sound of being needed for interviews in various places. “I’m quite pleased by the passion the fans have shown for me but believe me, I’m not the fan - I find the person out. It would be an extremely nice position to be in. Illuminate. We’d like to think, not to get a player based on Newcastle’s team when he does take charge. “My head coach says, ’Listen, it’s just too much to cry,’ Tony Wheeler told the BBC on Tuesday about the injury he played in the Northern League final at Paroles. “If the player is a part in the starting team, then I’m confident he would be put in there, but I also’ve finished telling players we thought we’d get the job. We had good options to go in the squad. “We had, obviously, the people that were always willing for us to get the best player we could because as coaches we felt we had the best players, and we felt that no one had a choice. “And the only thing [I] was just going on to say was that we were aware beforehand that it was an unusual situation for us to hook him up with the potential of [Meagan] Forster. “We like his game and he likes ours, and if he’s a good striker, he can go out there and play both for the team and for ourselves. “Stephen [Biggish] Jackson, who’s worked with here, obviously respects me with great respect and likes his opinions, but I think [Tony Wheeler] can say that it’s satisfactory. “We’ve got lots of many, and them combine them with Van Ferdinand. All three people we’ve been approached to, they need to be taken. “Could something be done to them? Not be done.” United are certainly prepared to lose their star man for Tuesday’s Rovers clash, with Borby to take charge of their trip to Dunferm on Saturday. Although they have won five of their six games, they sit sixth in the Northern League, which includes the Parkhead and the two assistant charges but on Tuesday their only goal of the season is worthy of their reward. United have more goals than Palace, scoring both goals this season despite the best form of Johnson as new coach Winger Darius Boyd joined Derby after replacing Marlon Wickham in June and recently, he stayed for a year on loan until he joined Newcastle in defiance of Tony Wheeler. Tom Williams, a future United captain, defender and striker, told BBC One Sports: “Being in Rovers is great for them. “They develop their chemistry very well in the dugout and they try to stay in good relationships together. “They’re unabashed about that you can play with both of them now, so they’ve stayed loyal.” Brendan Rodgers, the current Newcastle United manager after Wheeler took over from Mick Jagger, and succeeded David Moyes, as the most paid Premier League manager in 1997 while the Scot was the highest paid Everton manager. Leicester manager Leicester Leicester is yet to make terms with the joys of the job and has become so indebted he offered to forgive Newcastle United’s owners for the cost of lost profits after buying a stake in the club earlier this year. |endoftext| Sometimes, I enjoy having colleagues “embarrassed” to compete in sports, especially it comes to being a one-richen employee’s career. I can’t own the jobs of the colleagues that pressure us from Chinese and Australian competitive athletes and even to compete in the United States. They seem arrogant |endoftext|
@@
Figure 27: Sample from MDLM and $T=4096$ diffusion steps. |endoftext| token is bolded to emphasize the boundary between paragraphs.
@@
|endoftext| a component to renovating the park, but no one wants to involve others in it before facing a backlash. “We are going to stand up and try to make a real statement on this,” Thomas told reporters at a rally Tuesday night. “I would much rather have that there be a consensus not just to make a decision, but to have a very real discussion, and to begin to build a movement.” The idea behind the proposal is a call to the public, council and the youth advisory board, which was approved by residents on the first day of the meeting. Thomas said the message will be important to Hamilton youth. “It’s going to be very much about a number of people,” he said. “There will be arguments on both sides. It is not going to be all about the type of person that you are or the committee member that you are with a particular number of people.” Thomas said that the plan is really about an initiative to help the community do both of those things. “This isn’t going to be a positive thing, as it should be, but it’s going to put people in position to make sure that our focus is not just about how to deal with poverty …and with homelessness. Our focus is on keeping the people in our community affordable and to create opportunities for their families. “We have —believe it or not —we have increased community involvement in our city,” Thomas said. “I think it’s obvious from all of the issues now that we’ll consider resource-generating development and how to devribute those resources.” The renovated park is one of the 14 parks identified in the city’s 2008 map of the park’s 2-acre site, which was opened up for a community arts festival. City staff started coming up with the project in late December, but the public has yet to come to terms. For their part, residents sign the petition only to receive two differing opinions. Although there are concerns that it would be an uphill battle that requires council approval, officials said Tuesday that the process could eventually take shape, with councillors likely to support it. Hamilton city hall spokesman John Little told reporters that the city would have to address a petition from the public if there is one on site. “We are not able to change the outcome because we are discussing it internally, so what we have has not made an official decision,” Little said. Thomas said staff wants to hear how the public feels about the project, and it must be incorporated in such a prompt fashion that the city is looking at more permanent changes. “We implemented a 30-year plan, about 10 years ago, and we will continue it this year,” Thomas said. “We want to once again really emphasize community involvement. I mean, as we say we take this kind of step, we don’t want to make separate efforts to change different things,” he said. Any plans would be discussed with community boards, Conley said. Hamilton city councillor Steve Cowen said if residents sign the petition, “actions have to be done, knowing when these changes are needed and that we will have the right options to move forward.” Cowen said two renovations along with suggestions for revenue set by the Hulbert Company, which operates the Gerrard Elementary school theatre, will be included on city documents and website. They would set aside \$2 million more than the result of Washtenaw Park Plaza renovation plans, sparked by the QEDA expansion that renovated the former Washtenaw Historic Restaurant and Bakery on Gerrard. The company has also purchased Gerrard and turned it into a New American Food, Inc. restaurant, and brought the business before the city council in late 2015. Hamilton city council has been having this conversation for more than a year, but it usually happens only now. “It’s really not, ’We need a meeting soon,’ why we need to do this,” Conley said. “It’s just a forum where people can voice their views and inform council, and within the next year, hopefully council will do the right thing to do.” With John Elliott: 905-526-2431 mmelliott@thespec.com |endoftext| Written by Steven B. Ollander In the last five years, we have seen huge financial crises around the world, and not just economic ones. We’ve also seen rapid changes in economic systems, in central and peripheral countries. All these systems are in the hands of banks, but there are other forces that have also been used to dump more money into other countries |endoftext|